22 carat gold is mixed with 18 carat gold in the ratio 1:2 to get x carat gold . then the value of x is ?
Answer by Alan3354(69443)

22 carat gold is mixed with 18 carat gold in the ratio 1:2 to get x carat gold . then the value of x is ?
------------
22*1 + 18*2 = x*3
x = 58/3